OpenStudy (anonymous):

If h(x) = 3x − 1 and j(x) = −2x, solve h[j(2)] and select the correct answer below.

OpenStudy (anonymous):

−11 11 −13 13

OpenStudy (jdoe0001):

hmm, what's J(2) anyway?

OpenStudy (anonymous):

@zepdrix @mathmate

OpenStudy (anonymous):

j(2)=j(x) x=2. @jdoe0001

OpenStudy (jdoe0001):

well,, j(x) = -2x so if x = 2 that means j(2), then j(2) = ?

OpenStudy (anonymous):

-4

OpenStudy (jdoe0001):

yeap so now we know that j(2) = -4 thus \(\bf h({\color{brown}{ j(2)}}\iff h({\color{brown}{ -4}}=-3({\color{brown}{ -4}})-1\)

OpenStudy (anonymous):

okay so -4=11

OpenStudy (jdoe0001):

hmm \(\bf h({\color{brown}{ j(2)}})\iff h({\color{brown}{ -4}})=-3({\color{brown}{ -4}})-1 \\ \quad \\ h(j(2))=-12-1\implies h(j(2))=-13\)

OpenStudy (anonymous):

so c, -13? Thanks for the help. Can you help with one more?

OpenStudy (jdoe0001):

hmm actually have a typo, is 13 no -13 but anyhow \(\bf h({\color{brown}{ j(2)}})\iff h({\color{brown}{ -4}})=3({\color{brown}{ -4}})-1 \\ \quad \\ h(j(2))=-12-1\implies h(j(2))=-13\)

OpenStudy (jdoe0001):

sure, post anew, thus more eyes

OpenStudy (anonymous):

so 13

OpenStudy (anonymous):

not -13

OpenStudy (jdoe0001):

3* -4 = -12 -12 - 1 = -13

OpenStudy (anonymous):

so the answer is -13?

OpenStudy (jdoe0001):

well... the answer is \(\bf h({\color{brown}{ j(2)}})\iff h({\color{brown}{ -4}})=3({\color{brown}{ -4}})-1 \\ \quad \\ h(j(2))=-12-1\implies h(j(2))=-13\) :)

OpenStudy (anonymous):

but its multiple choice, so h(j(2))=-13

OpenStudy (anonymous):

thanks!

OpenStudy (jdoe0001):

yw